OpenStudy (anonymous):

f(x)=(x^2)(e^5x) find f"(x).

OpenStudy (hoblos):

f(x) = u(x) v(x) f'(x) = u'(x)v(x) + u(x)v'(x) use this formula to get the answer

OpenStudy (anonymous):

o trust me we have tried lol

OpenStudy (anonymous):

you will have to use the quotient rule and that's what hobios has shown....

OpenStudy (anonymous):

whoops product rule!!!

OpenStudy (hoblos):

f'(x) = (2x)(e^5x) + 5(x^2)(e^5x) f''(x) = 2e^5x + 10x(e^5x) + 10x(e^5x) + 25(x^2)(e^5x)

OpenStudy (anonymous):

I just did it I dodn't have any big issues. Just use the produle rule 1 time for the f'(x) and twice for the f''(x)...shouldn't be nothing tricky

OpenStudy (anonymous):

just remember the derivative of e^x = e^x * x'

OpenStudy (hoblos):

f''(x) = (e^5x)(2+20x + 25x^2)

OpenStudy (anonymous):

yup thats it. thanks.

OpenStudy (hoblos):

its my pleasure XD

OpenStudy (anonymous):

can you show me how you got f''(x) = 2e^5x + 10x(e^5x) + 10x(e^5x) + 25(x^2)(e^5x)

OpenStudy (anonymous):

i know u used the product rule but i dont see it

OpenStudy (hoblos):

use product rule for each part alone then add them [(2x)(e^5x)]' = 2e^5x + 10x(e^5x) [5(x^2)(e^5x)]' = 10x(e^5x) + 25(x^2)(e^5x)

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4f4f9a9de4b019d0ebaed1ed HELP?